OpenStudy (anonymous):

An overnight delivery service accepts packages of up to 165 inches in length and girth combined. (Girth is the distance around the package.) A package has a fixed girth of 53 inches. Determine (in terms of an inequality) those lengths for which a package is acceptable. Write your answer in the form of an inequality.

OpenStudy (anonymous):

i like how like 5 people are viewing and no one can answer it. haha

OpenStudy (anonymous):

This is easy math problem but written badly

OpenStudy (anonymous):

i agree.

OpenStudy (lalaly):

\[L+g \le 165\]\[g=53\]\[L+53 \le 165\]\[L \le 112\]

OpenStudy (anonymous):

Thank you :D

OpenStudy (lalaly):

your welcome:)